# Robo3R: Enhancing Robotic Manipulation with Accurate Feed-Forward 3D Reconstruction

### v1

@@ II Related Work
3D representation for manipulation. Compared to RGB images, 3D representation provides better physical grounding and are therefore widely used in robotic manipulation. Several studies have employed 3D representation as model inputs to enable efficient and generalizable policy learning [49, 43, 9, 26, 48, 39, 12, 10, 11, 42, 22]. It has also been shown that 3D representation is more suitable than RGB images for sim-to-real transfer of manipulation policies [23, 46, 25, 20, 47]. Moreover, grasp synthesis models typically rely on 3D representation to model the scene [7, 52, 37, 32, 17, 41], thereby enabling the prediction of grasp poses grounded in 3D geometry. Collision-free motion planning in real-world environments also requires 3D representation [28, 3, 44, 14], as explicit 3D data provide superior spatial and geometric information for obstacle avoidance. Recently, several studies leverage implicit 3D representation to enhance the performance of manipulation policies [16, 24, 51, 19]. We evaluate the quality and usability of the 3D representation reconstructed by Robo3R in downstream applications such as imitation learning, sim-to-real transfer, grasp synthesis, and collision-free motion planning.
@@ III-B3 Prediction Heads
Masked point head. Over-smoothing is a persistent issue in dense prediction, which could result in blurry edges and loss of detail in point clouds. To address this problem, we decouple dense point prediction into depth, image coordinate, and mask components, as illustrated in Fig. 3. The mask is used to separate the depth and image coordinates corresponding to different scene parts. Subsequently, the depth values are unprojected into 3D points using the image coordinates, and the points of all parts are aggregated to form the whole scene. Specifically, the masked point head consists of a five-layer transformer decoder, a depth head, a ray head, and a mask head. Patch-wise features produced by the transformer backbone are processed by the transformer decoder. The outputs are then decoded into depth, image coordinates, and masks using the MLP heads followed by pixel shuffle operations.
@@
Extrinsic estimation module and keypoint head. Although Robo3R directly predicts the camera extrinsics, which are part of the similarity transformation, we additionally develop a more accurate and robust extrinsic estimation module, as illustrated in Fig. 4. The results obtained from this module can further refine the similarity transformation. We predefine keypoints on the robot kinematic chains and use forward kinematics to obtain their 3D positions. Robo3R predicts the pixel coordinates of these keypoints with a keypoint head, and the camera extrinsics are subsequently estimated by solving a PnP problem. In the keypoint head, patch features are first processed by a transformer decoder, then decoded by an MLP followed by a pixel shuffle operation to generate a keypoint heatmap $M_{kp}\in\mathbb{R}^{H\times W\times N_{kp}}$, where $N_{kp}$ is the number of keypoints. Finally, we employ a differentiable Soft-Argmax operator to extract 2D keypoint coordinates $C_{kp}\in\mathbb{R}^{N_{kp}\times 2}$ from the heatmaps with sub-pixel accuracy.
@@ IV-A1 Quantitative Performance
Benchmark. We constructed a benchmark for 3D reconstruction in robotic manipulation scenarios by rendering a photorealistic and diverse dataset, using objects, textures, and environment maps that are different from those used in training. The test set contains 2,000 scenes and 80,000 frames. The evaluation consists of two aspects: point map estimation and relative camera pose estimation. The metrics for point map estimation include scale-invariant point error, normal error, and scale error. For relative camera pose estimation, the metrics include relative translation error (RTE), relative rotation error (RRE), relative translation accuracy (RTA) at a given threshold (e.g., RTA@0.03 for 0.03 meters), and relative rotation accuracy (RRA). As robots typically perceive the environment from sparse views, we evaluate point map estimation under monocular and binocular settings, and assess relative camera pose estimation under the binocular setting.
Baselines. We select four leading feed-forward 3D reconstruction models as baselines: VGGT [30], $\pi^{3}$ [36], MapAnything (MA) [13], and DepthAnything3 (DA3) [18]. Among them, MA and DA3 perform metric-scale reconstruction.
Results. We evaluate Robo3R against strong baselines on the benchmark. Tab. I presents the quantitative results for point map estimation. Our method significantly outperforms all baselines across all metrics in both monocular and binocular settings. In the monocular case, Robo3R achieves a point error of 0.006, representing an order of magnitude improvement over the second-best method, $\pi^{3}$. Unlike the other models, which suffer from severe scale ambiguity (scale errors >0.46>0.46), Robo3R effectively recovers metric geometry. This superiority persists in the binocular setting, where Robo3R further reduces errors, demonstrating robustness in 3D reconstruction for challenging manipulation scenarios. As shown in Tab. II, our method demonstrates exceptional precision in relative pose estimation. Robo3R achieves an RTE of 0.014 and an RRE of 0.013, which are approximately $8\times$ and $5\times$ lower than those of the best baseline, $\pi^{3}$, respectively. Furthermore, the high accuracy rates (RTA@0.03 of 0.951) confirm that Robo3R consistently provides reliable relative camera pose.
TABLE I: Results on point map estimation. We report the scale-invariant point error, normal error, and scale error for each method.
Method Point Err. $\downarrow$ Normal Err. $\downarrow$ Scale Err. $\downarrow$
Monocular VGGT 0.126 0.231 0.663
$\pi^{3}$ 0.061 0.140 0.497
MA 0.078 0.143 0.467
DA3 0.075 0.138 0.506

### v2

@@ II Related Work
3D representation for manipulation. Compared to RGB images, 3D representation provides better physical grounding and are therefore widely used in robotic manipulation. Several studies have employed 3D representation as model inputs to enable efficient and generalizable policy learning [51, 44, 9, 27, 50, 40, 12, 11, 10, 43, 22, 24]. It has also been shown that 3D representation is more suitable than RGB images for sim-to-real transfer of manipulation policies [23, 48, 26, 20, 49]. Moreover, grasp synthesis models typically rely on 3D representation to model the scene [7, 54, 38, 35, 17, 42, 47], thereby enabling the prediction of grasp poses grounded in 3D geometry. Collision-free motion planning in real-world environments also requires 3D representation [29, 3, 45, 14], as explicit 3D data provide superior spatial and geometric information for obstacle avoidance. Recently, several studies leverage implicit 3D representation to enhance the performance of manipulation policies [16, 25, 53, 19]. We evaluate the quality and usability of the 3D representation reconstructed by Robo3R in downstream applications such as imitation learning, sim-to-real transfer, grasp synthesis, and collision-free motion planning.
@@ III-B3 Prediction Heads
Masked point head. Over-smoothing is a persistent issue in dense prediction, which could result in blurry edges and loss of detail in point clouds. To address this problem, we decouple dense point prediction into depth, image coordinate, and mask components, as illustrated in Fig. 3. The masked point head consists of three branches dedicated to the robot, objects, and background. Each branch contains a depth head, a ray head, and a mask head, and produces a point map by unprojecting the depth along the corresponding ray directions. The robot, object, and background masks then extract region-specific points from their respective point maps, which are aggregated to reconstruct the full scene point cloud. Specifically, the masked point head consists of a five-layer transformer decoder, a depth head, a ray head, and a mask head. Patch-wise features produced by the transformer backbone are processed by the transformer decoder. The outputs are then decoded into depth, image coordinates, and masks using the MLP heads followed by pixel shuffle operations.
@@
Extrinsic estimation module and keypoint head. Although Robo3R directly predicts the camera extrinsics, which are part of the similarity transformation, we additionally develop a more accurate and robust extrinsic estimation module, as illustrated in Fig. 4. The results obtained from this module can further refine the similarity transformation. We compute 3D keypoints as the frame origin of each link by applying forward kinematics to the robot’s joint positions. During data synthesis, these keypoints are projected onto the image plane via camera intrinsics and extrinsics to obtain 2D keypoints, each rendered as a Gaussian blob to produce a per-keypoint heatmap. During inference, Soft-Argmax is applied to heatmaps to extract 2D keypoints as a weighted average of pixel coordinates, yielding floating-point values that enable sub-pixel accuracy. The 3D keypoints and per-keypoint heatmaps are index-aligned, enabling matching between 3D and 2D keypoints. Robo3R predicts the pixel coordinates of these keypoints with a keypoint head, and the camera extrinsics are subsequently estimated by solving a PnP problem. In the keypoint head, patch features are first processed by a transformer decoder, then decoded by an MLP followed by a pixel shuffle operation to generate a keypoint heatmap $M_{kp}\in\mathbb{R}^{H\times W\times N_{kp}}$, where $N_{kp}$ is the number of keypoints. Finally, we employ a differentiable Soft-Argmax operator to extract 2D keypoint coordinates $C_{kp}\in\mathbb{R}^{N_{kp}\times 2}$ from the heatmaps with sub-pixel accuracy.
@@ IV-A1 Quantitative Performance
Benchmark. We constructed a benchmark for 3D reconstruction in robotic manipulation scenarios by rendering a photorealistic and diverse dataset, using objects, textures, and environment maps that are different from those used in training. The test set contains 2,000 scenes and 80,000 frames. The evaluation consists of two aspects: point map estimation and relative camera pose estimation. The metrics for point map estimation include scale-invariant point error and scale error. For relative camera pose estimation, the metrics include relative translation error (RTE), relative rotation error (RRE), relative translation accuracy (RTA) at a given threshold (e.g., RTA@0.03 for 0.03 meters), and relative rotation accuracy (RRA). As robots typically perceive the environment from sparse views, we evaluate point map estimation under monocular and binocular settings, and assess relative camera pose estimation under the binocular setting.
Baselines. We select four leading feed-forward 3D reconstruction models as baselines: VGGT [31], $\pi^{3}$ [37], DepthAnything3 (DA3) [18], and MapAnything (MA) [13]. Among them, MA and DA3 perform metric-scale reconstruction. Additionally, we consider MapAnything fine-tuned on the Robo3R-4M dataset (MA-FT) as a baseline.
Results. We evaluate Robo3R against strong baselines on the benchmark. Tab. I presents the quantitative results for point map estimation. Our method significantly outperforms all baselines across all metrics in both monocular and binocular settings. In the monocular case, Robo3R achieves a point error of 0.006, representing an order of magnitude improvement over the second-best method, $\pi^{3}$. Unlike the other models, which suffer from severe scale ambiguity (scale errors >0.46>0.46), Robo3R effectively recovers metric geometry. This superiority persists in the binocular setting, where Robo3R further reduces errors, demonstrating robustness in 3D reconstruction for challenging manipulation scenarios. Additionally, Robo3R consistently outperforms the fine-tuned MapAnything model (MA-FT). This highlighting the benefits of our proposed model architecture. As shown in Tab. II, our method demonstrates exceptional precision in relative pose estimation. Robo3R achieves an RTE of 0.014 and an RRE of 0.013, which are approximately $8\times$ and $5\times$ lower than those of the best baseline, $\pi^{3}$, respectively. Furthermore, the high accuracy rates (RTA@0.03 of 0.951) confirm that Robo3R consistently provides reliable relative camera pose.
TABLE I: Results on point map estimation. We report the scale-invariant point error and scale error for each method.
Method Point Err. $\downarrow$ Scale Err. $\downarrow$
Monocular VGGT 0.126 0.663
$\pi^{3}$ 0.061 0.497
DA3 0.075 0.506
MA 0.078 0.467
MA-FT 0.010 0.010
@@ IV-C Effectiveness of Design Choices
We provide qualitative comparisons for the masked point head, since its advantage in producing finer-grained geometry is not well captured by conventional benchmarks. Specifically, we train a variant that retains the point decoder architecture but removes the masked design (w/o MPH). As shown in Fig. 8, the masked point head enables the model to predict clean points on structures as fine as 1.5 mm, whereas the baseline fails to recover such geometry.